# An Inversion-based Measure of Memorization for Diffusion Models

### v2

@@ 8.2 Detection
Experiment details. On CIFAR-10, the training loss metrics are calculated on the average loss of 16 random Gaussian noise and 50 timesteps uniformly sampled within the range [1, 1000]. Following van den Burg et al. [54]’s setting, $M_{\mathrm{LOO}}$ is estimated using a 10-fold cross-validation. We only conduct the cross-validation once.
On LAION, the training loss metrics use 32 random Gaussian noise and 50 timesteps uniformly sampled within the range [1, 1000]. We implement Wen et al. [61] following their best performing settings.
@@ 8.3 Collate InvMM with nearest neighbor test
Epoch ×$\times 2$ 0.059(9) N/A(0) N/A(0) N/A(0) - -
Epoch ×$\times 3$ 0.326(120) N/A(0) 0.000(1) N/A(0) - -
Epoch ×$\times 4$ 0.567(181) 0.000(3) 0.000(4) N/A(0) - -
Duplicate ×$\times 5$ 0.182(65) - 0.000(2) - -
Duplicate ×$\times 10$ 0.852(237) - 0.192(186) - -
Duplicate ×$\times 15$ 0.937(245) - 0.401(323) - -
Duplicate ×$\times 20$ 0.819(231) - 0.434(375) - -
@@
Under the setting in Sec. 5.3, InvMM achieves an IoU of 0.817 on CIFAR-10 and 1.0 on LAION. The results on CelebAHQ and FFHQ are summarized in Tab. 3. 10k random samples are used to obtain the results. A random sample of cosine similarity larger than 0.5 with any training sample is considered a replication. InvMM presents consistence with the results of nearest neighbor test, indicating its potential to expose risk of training image leakage, especially when a large number of training images are prone to leakage. The 10k samples is not an adequate sampling of the large latent space (64$\times$64$\times$3), the performance can be refined with a larger scale of evaluation.
@@ 8.4 Comparison to membership
Experiment details. Figure 9 is evaluated with LDM trained on FFHQ-6k. FFHQ-600 is used as the member set and CelebAHQ-250 is the hold-out set. The statistic variant of SecMI is implemented following their official settings, with $t_{\mathrm{SEC}}=100$ and $k=10$.
@@
Results. Figure 11 provides quantitative evidence that different samples could be replicated with different levels of training loss. The replication loss is calculated over the inverted latent noise distribution. It shows that although holdout samples have a larger training loss over the $\mathcal{N}(0,I)$, supporting effective membership inference, it is still possible to replicate some of them by reducing their training loss to a level higher than some of the members.
@@ 9.2 Experiment setup
We utilize SD v1.4 for evaluation, which uses CLIP [39] to encode input prompts and generate high-resolution $512\times 512$ images. The text encoders by default accepts a maximum length of 77 tokens, in which the first and last tokens are padded tokens indicating the start and end of a prompt. The rest 75 tokens are all optimized in our experiments. During optimization, the parameters $\bm{\theta}$ of diffusion models are fixed. We optimize for 500 iterations with a constant learning rate of 0.2. $\bm{\phi}$ is initialized to $\mathbf{0}$ at the begining of optimization.
@@ 9.3 Influence of temperature
Denoising error. We first analyze the denoising error of the inverted prompt distribution when optimized with different temperatures. We consider 4 settings for the temperature to be either 0.1, 0.5, 1 or 2, as well as whether to discretize the smoothed tokens $\tilde{\bm{\omega}}$ to $\bm{\omega}$. After optimization, random prompts and noises can be drawn from $q_{\bm{\phi}}(\bm{\omega}|\bm{x}_{0})$ and $\mathcal{N}(\mathbf{0},\mathbf{I})$. For each image, we randomly sample 10 prompts and 10 noises for each sampled prompt, resulting in 100 prompt-noise pairs. The denoising error is estimated using 50 timesteps uniformly sampled within the range $[1,1000]$ and averaged over the 100 prompt-noise pairs.
@@
9.4 Alternative prompt tuning methods
Figure 18: Inversion results of AutoPrompt and PEZ. Within each block, left: optimize 75 tokens, right: optimize tokens equal to training captions.
We also compare to several existing discrete prompt tuning methods, to see if they work for training image inversion. Figure 16 has shown that GBDA [20] fails ($\tau=1.0$ with discretization). We further study the effectiveness of AutoPrompt [47] and PEZ [60]. AutoPrompt and PEZ are implemented following their default settings. We evaluate with 4 heavily memorized images (almost 100% sensitivity on their captions). In addition to optimizing all the 75 tokens, we also set the number of optimizable tokens the same as the corresponding training captions (shorter), to reduce the difficulty of inversion. Figure 18 shows the inversion results. AutoPrompt and PEZ both fail to reconstruct the original training images.
Conclusion. Existing discrete prompt tuning methods cannot effectively work to eidetically reconstruct training images. We temporally resort to relaxed prompt tuning.
@@ 9.7 Adaptive algorithm
9.8 Full results
Based on the previous ablation study, we investigate memorization with the appropriate choice of each factor. In addition to SD v1.4, we include SD v2.1 for comparison. When not otherwise specified, experiment settings are the same as ablation study.
Figure 22: Number of training data replication in the confirmed group, sorted in descending order.
We first conduct a complete check of generation results. For each training image, 10 random prompts are sampled from $q_{\bm{\phi}}(\bm{\omega}|\bm{x}_{0})$ and for each sampled prompt, 100 Gaussian noises are sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. With an array of CFG scales, each prompt results in 700 randomly generated images. We manually count the number of generated images that present an obvious replication of the corresponding training image and report the maximum number of replications among the 10 prompts.
Figure 22 shows the results of the confirmed set for SD v1.4. There are 61 images being 100% replicated and the rest 15 images are replicated at least 5 times and at most 96 times. In the suspicious set, there are 3 newly discovered images that are replicated 53, 100 and 87 times. This demonstrates the advantage of our inversion-based analysis method, because some images cannot be directly generated by their training captions but do present a relatively large extent of memorization, compared to images in the normal set, where we found no replication.
Figure 23: Inversion results on SD v2.1 (Fixed noise distribution). The first column shows the corresponding training images.
For SD v2.1, we have not observed any replication but very similar generations, shown in Fig. 23. This is probably because SD v2.1 deduplicated the training set, and thus the true likelihood of previous duplicate samples in the training set is significantly reduced.
The counts demonstrate that training images are memorized to varied extent: For some images, training data leakage is certain to happen, for some images, it happens by chance and for the other images, it hardly happens. We also examine the images that present partial replication in our experiments by generating using their training captions and find that indeed, they are not memorized as severely as those that are completely memorized. It thus motivates us to find a measure to tell the difference.
@@ 10 More generation results
Figures 24, 25, 26, 27, 28, 29, 30 and 31 show more generated images using inverted noise vectors (and prompts).

### v3

@@ 8.3 Detection
Experiment details. On CIFAR-10, the training loss metrics are calculated on the average loss of 16 random Gaussian noise and 50 timesteps uniformly sampled within the range [1, 1000]. Following van den Burg et al. [55]’s setting, $M^{\mathrm{LOO}}$ is estimated using a 10-fold cross-validation.
On LAION, the training loss metrics use 32 random Gaussian noise and 50 timesteps uniformly sampled within the range [1, 1000]. We implement Wen et al. [61] and Ren et al. [42] following their best performing settings. For GCG [64] attack, the number of optimizable tokens is 20. Each token is initialized to the special token $<|endoftext|>$. GCG is ran for 500 steps with a batch size of 128 and a top-k of 256. During optimization, the noise distribution is fixed to the standard Gaussian. A minibatch of 16 random noises is used to calculate the $\bm{x}_{0}$-loss every iteration. After each update, 8 random samples are generated. If any of them has a similarity with the target image no less than 0.5, the optimization process will stop.
@@
Figure 13: Additional examples of adversarial prompts found by GCG.
Adversarial prompts. Figure 13 shows additional examples of adversarial prompts found by GCG for images from the suspicious set, which have lower InvMM scores.
@@ 8.4 Collate InvMM with nearest neighbor test
Epoch ×2$\times$ 2× 2 0.059(9) N/A(0) N/A(0) N/A(0) - -
Epoch ×3$\times$ 3× 3 0.326(120) N/A(0) 0.000(1) N/A(0) - -
Epoch ×4$\times$ 4× 4 0.567(181) 0.000(3) 0.000(4) N/A(0) - -
Duplicate ×5$\times$ 5× 5 - 0.182(65) - 0.000(2) - -
Duplicate ×10$\times$ 10× 10 - 0.852(237) - 0.192(186) - -
Duplicate ×15$\times$ 15× 15 - 0.937(245) - 0.401(323) - -
Duplicate ×20$\times$ 20× 20 - 0.819(231) - 0.434(375) - -
@@
Under the setting in Sec. 5.4, InvMM achieves an IoU of 0.817 on CIFAR-10 and 1.0 on LAION. The results on CelebAHQ and FFHQ are summarized in Tab. 3. 10k random samples are used to obtain the results. A random sample of cosine similarity larger than 0.5 with any training sample is considered a replication. InvMM presents consistence with the results of nearest neighbor test, indicating its potential to expose risk of training image leakage, especially when a large number of training images are prone to leakage. The 10k samples are not an adequate sampling of the large latent space (64$\times$64$\times$3), the performance can be refined with a larger scale of evaluation.
@@ 8.5 Comparison to membership
Experiment details. Figure 10 is evaluated with LDM trained on FFHQ-6k. FFHQ-600 is used as the member set and another 600 images not contained in FFHQ-6k constitute the hold-out set. The statistic variant of SecMI is implemented following their official settings, with $t_{\mathrm{SEC}}=100$ and $k=10$.
@@
Results. Figure 14 provides quantitative evidence that different samples could be replicated with different levels of training loss. The replication loss is calculated over the inverted latent noise distribution. It shows that although holdout samples have a larger training loss over the $\mathcal{N}(\mathbf{0},\mathbf{I})$, supporting effective membership inference, it is still possible to replicate some of them by reducing their training loss to a level higher than some of the members.
@@ 9.2 Experiment setup
We utilize SD v1.4 for evaluation, which uses CLIP [40] to encode input prompts and generate high-resolution $\times 512$ 512 images. The text encoders by default accepts a maximum length of 77 tokens, in which the first and last tokens are padded tokens indicating the start and end of a prompt. The rest 75 tokens are all optimized in our experiments. During optimization, the parameters $\bm{\theta}$ of diffusion models are fixed. We optimize for 500 iterations with a constant learning rate of 0.2. $\bm{\phi}$ is initialized to $\mathbf{0}$ at the begining of optimization.
@@ 9.3 Influence of temperature
Denoising error. We first analyze the denoising error of the inverted prompt distribution when optimized with different temperatures. We consider 4 settings for the temperature to be either 0.1, 0.5, 1 or 2, as well as whether to discretize the smoothed tokens $\tilde{\bm{\omega}}$ to $\bm{\omega}$. After optimization, random prompts and noises can be drawn from $q_{\bm{\phi}}(\bm{\omega}|\bm{x}_{0})$ and $\mathcal{N}(\mathbf{0},\mathbf{I})$. For each image, we randomly sample 10 prompts and 10 noises for each sampled prompt, resulting in 100 prompt-noise pairs. The denoising error is estimated using 50 timesteps uniformly sampled within the range [1,1000][1,1000][ 1 , 1000 ] and averaged over the 100 prompt-noise pairs.
@@ 10 More generation results
Figures 24, 25, 26, 27, 28, 29, 30, 31, 32 and 33 show more generated images using inverted noise vectors (and prompts).
@@
Figure 30: Random samples of SD v3.5 inversion on the confirmed subset. The first column shows the corresponding training images.
Figure 31: Random samples of SD v3.5 inversion on the suspicous subset. The first column shows the corresponding training images.